OpenStudy (anonymous):

if 6 wooden disks are cut from a 36 centimeter by 24 centimeter rectangular block of wood how many square centimeters of wood will not be used

OpenStudy (ddcamp):

 \[Area\ of \ 1\ circle= 36\pi\] \[Area\ of\ all\ 6\ circles\ = 216\pi\] \[Area\ of\ rectangle = 36 * 24 = 864\] \[Area\ left\ over = 864-216\pi = 185.416 \]